OpenStudy (anonymous):

for this sequence -1 -7 -13 -19 -25 would this function rule be correct? f(x)=x-6

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

I think it is because the graph is going by -6 every time or well it is taking away six every time.

OpenStudy (anonymous):

ehh... in reality it depends what you call the first value of the sequence :P. is that x= 0 or x=1.

OpenStudy (anonymous):

it's f(n) = -1 -6n where n=0 is the first element.

OpenStudy (anonymous):

you can change it if you want n=1 to be the first element

OpenStudy (anonymous):

what you have defined is called a "recursion" meaning the next term is 6 less than the last one. it would be written as \[a_n=a_{n-1}-6\] but i do't think that is what is asked for

OpenStudy (anonymous):

Ha i dont know.. the question just says write a function rule for that sequence :P its supposed to be along the lines of f(x) (stuff here)

OpenStudy (anonymous):

o_o then refer to me above :P

OpenStudy (anonymous):

since you are decreasing by 6 each time, then you have a line with slope -6

OpenStudy (anonymous):

Ha so i did it right !?

OpenStudy (anonymous):

no you have the right idea, but i think you need something like \[f(x)=-6x+5\]

OpenStudy (anonymous):

if you are looking for a function. if you are looking for the rule for a sequence then it would be \[a_n=a_{n-1}-6, a_1=-1\]

OpenStudy (anonymous):

Oh.. well atleast i had the right idea :)